# Rank jumps and Multisections of elliptic fibrations on K3 surfaces

### v1

@@ 3.2.1. Automorphisms of K3 surfaces in $\mathcal{L}_{d}$
Let $X$ be a general K3 surface in a family $\mathcal{L}_{d}$. Fix the basis $\{u_{1},u_{2},b_{3}\}$ of $\Lambda_{d}$ considered in proof of Lemma 3.8. The elliptic involution on each fiber of $\pi$ extends to an automorphism on $X$ that acts on the chosen basis of $\Lambda_{d}$ as follows:
@@
If $d\geq 2$, the translation by the section $S_{1}$ is an automorphism of infinite order of the K3 surface, which acts on the basis $\{F,\mathcal{O},b_{3}\}$ as follows:
@@
If $d=1$, the previous matrix is still associated with an isometry of the lattice $\Lambda_{d}$, but this isometry cannot be associated with an automorphism since it maps $F-b_{3}$ to $-F+b_{3}$. The former is an effective class corresponding to the component of $I_{2}$-fiber meeting the zero section. The latter is not effective since $(-F+b_{3})\mathcal{O}=-1$.
@@
If $X$ is general in its family, then it is $\mathrm{Aut}$-general, i.e., its automorphism group contains only symplectic automorphisms of infinite order (e.g. $T_{1}$) and non-symplectic automorphisms of order 2 (e.g. $\varepsilon$). In particular, this implies that $X$ does not admit isotrivial elliptic fibrations with $j$-invariant of the generic fiber equal to 00 or $1728$. Indeed, otherwise the complex multiplication on each fiber would induce an automorphism on the surface which is non–symplectic and has order 3 or 4, respectively. If an elliptic fibration is isotrivial with a constant $j$-invariant not equal to 0 or 1728, then the reducible fibers are forced to be of type $I_{0}^{*}$, and in particular with non reduced components. By Lemma 2.13, these isotrivial fibrations cannot appear if the Picard number of $X$ is at most 3, i.e. for the surfaces considered in this paper.
@@ 3.3. Smooth rational bisections on elliptic K3 surfaces
Let $X$ be a general K3 surface in $\mathcal{F}_{\rm{rat}}^{x}$ and recall that $x\geq 0$ since it is the intersection number of two irreducible distinct curves. The classes $F$, of the fiber of the fibration, $\mathcal{O}$, of the zero section of the fibration, and $B$, of the rational bisection, are contained in the Picard lattice of $X$. Let us denote $\Phi_{x}:=\left[\begin{array}[]{rrr}0&1&2\\ 1&-2&x\\ 2&x&-2\end{array}\right]$ the lattice $\langle F,\mathcal{O},B\rangle$. It is primitively embedded in $\operatorname{Pic}(X)$, since it does not admit even finite index overlattices in which $\langle F,\mathcal{O}\rangle$ and $\langle B\rangle$ are primitively embedded. Hence, we proved that the family $\mathcal{F}_{\rm{rat}}^{x}$ is the family of the $\Phi_{x}$-polarized K3 surfaces.
To compare the families $\mathcal{F}_{\rm{rat}}^{x}$ and $\mathcal{L}_{2x+5}$, it suffices to show that the lattices $\Phi_{x}$ and $\Lambda_{2x+5}$ are isometric. So we consider a different basis for $\operatorname{Pic}(X)$: $\{F,F+\mathcal{O},(-4-x)F-2\mathcal{O}+B\}$. The intersection form on this basis is $U\oplus\langle-2(2x+5)\rangle$. This proves that $\operatorname{Pic}(X)$ is isometric to $\Lambda_{d}$ if $d=2x+5$, with $x\geq 0$.∎
@@
If $\operatorname{NS}(X)\simeq\Lambda_{3}$ (i.e. $d=3$), then the class $B=3F+2\mathcal{O}+b_{3}$ as in (3.2) is still a $(-2)$-class with intersection 2 with $F$, but it does not correspond to an irreducible curve, indeed $B\mathcal{O}=-1<0$ and $\mathcal{O}$ is an irreducible curve. Indeed, the class $B$ is not an irreducible bisection, but the sum of two sections, namely $S_{1}=3F+\mathcal{O}+b_{3}$ and $\mathcal{O}$.
@@
The cover involution of $f:X\rightarrow\mathbb{P}^{2}$ is the involution $\varepsilon\circ T_{1}$ and preserves the class of the bisection $B$.
@@ 3.4. Smooth genus 1 bisections on elliptic K3 surfaces
The proof is analogous to the one of the case of rational bisection (see Proposition 3.13), by observing that the classes $F,\ \mathcal{O},\ B$ have the following intersection properties $F^{2}=B^{2}=0$, $F\mathcal{O}=1$, $FB=2$ and $B\mathcal{O}=x$ and that there is a base change which gives $u_{1}=F$, $u_{2}=F+\mathcal{O}$ and $b_{3}=-(x+4)F-2O+B$.∎
@@
The base change between $\{F,\mathcal{O},B\}$ and $\{u_{1},u_{2},b\}$ gives $B=b_{3}+(4+x)F+2\mathcal{O}=(2+x)u_{1}+2u_{2}+b_{3}$, and this allows one to compute the image of $B$ for the automorphisms $T_{1}$, $\varepsilon$ and their composition:
@@
The cover involution of $f:X\rightarrow\mathbb{P}^{1}\times\mathbb{P}^{1}$ is the involution $\varepsilon\circ T_{1}$. It preserves the class of the bisection $B$.
@@ 4. Higher degree multisections
We consider here multisections of degree higher than 2. The main ideas and the proofs are very similar to those in the previous section, where we considered multisections of degree 2. Indeed, as in that case, one can determine the 18-dimensional (resp. 17-dimensional) families of K3 surfaces with a genus 1 fibration (resp. elliptic fibration) admitting a smooth rational or genus 1 multisection of degree $m>2$ and for low values of $m$, namely $m=3,4,5$; moreover, one is also able to describe projective models of the generic K3 surface in the constructed families. The main difference with respect to the case of degree $m=2$ is that now there could exist genus 1 fibrations admitting smooth multisections of degree $m>2$ but none of them are rational or of genus 1.
@@
Let $X$ be a K3 surface with a genus 1 fibration and a multisection of degree $m$ that is rational (resp. of genus 1). Then $X$ is $\left[\begin{array}[]{cc}0&m\\ m&2c\end{array}\right]$-polarized with $c=-1$ (resp. $c=0$).
We observe that if $m=3$, and $X$ is a K3 surface admitting a genus 1 fibration with a trisection of low genus, $X$ can be polarized with three non-isometric lattices that are $\left[\begin{array}[]{cc}0&3\\ 3&2c\end{array}\right]$, with $c=-1,0,1$. They correspond to the cases in which the trisection has genus 0,1 or 2, respectively. When $m$ grows, more possibilities for the polarizing lattices appear, see [vG05, Proposition 3.7].
For low values of $m$, the previous Lemma gives nice geometric models of the K3 surfaces and the genus 1 fibration (see [vG05, Paragraph 5.2]): let $M$ be a smooth rational multisection of degree $m$ of the genus 1 fibration whose class of the fiber is $F$, then $H:=F+M$ is an ample divisor, $\varphi_{|H|}(X)\subset\mathbb{P}^{m}$ and there is a pencil of hyperplanes in $\mathbb{P}^{m}$ containing the degree $m-2$ curve $C:=\varphi_{|H|}(M)$.
@@
In all the cases above, the singular fibers of the fibration correspond to hyperplanes which cut on $\varphi_{|H|}(X)$ the curve $C$ and a singular curve. Since $M$ is a multisection of the fibration, we have a map $M\rightarrow\mathbb{P}^{1}$ to the basis of the fibration. It is branched over fibers cut out by hyperplanes whose intersection with $\varphi_{|H|}(X)$ consists of $C=\varphi_{|H|}(M)$ and a curve which intersects $C$ in at least one point with multiplicity higher than 1. For a general K3 surface, the hyperplanes corresponding to singular fibers and the ones corresponding to branch fibers for $M\rightarrow\mathbb{P}^{1}$ do not coincide and hence the multisection $M$ is generically a saliently ramified multisection.
@@
Similarly to the proofs of Proposition 3.13 and Lemma 3.19, it suffices to look for an isometry between the lattice $\langle F,\mathcal{O},M\rangle$ and a lattice $\Lambda_{d}$ for a certain integer $d$. Assuming $M\mathcal{O}=x\geq 0$, one obtains $d=m^{2}+xm+1$, if $M$ is rational, and $d=m^{2}+2mx$, if it is of genus 1.∎
@@
In Theorem 3.2, we find smooth bisections of low genus (0 or 1) on the K3 surfaces in all the families $\mathcal{L}_{d}$ with $d\geq 4$; the smooth multisections of low genus and degree $m>2$ appear for higher values of $d$. So we do not find smooth saliently ramified multisections of degree $m>2$ on elliptic fibrations that do not admit smooth saliently ramified bisections. We conclude that
@@ 5. Rank jumps for K3 surfaces in $\mathcal{L}_{d}$
Therefore, the only statement which has not yet been proved concerns the field of definition $k$, namely, to show that the sets $\mathcal{R}(X,\pi,k)$ (resp. $\mathcal{R}(X,\pi,l)$) are infinite if $d\equiv 3\mod 4$ (resp. $d\equiv 1\mod 4$). In particular will verify that $B$ is defined over $k$ if $d$ is odd, that it has infinitely many $k$-points if $d\equiv 3\mod 4$ and that it has infinitely many $l$-points for a certain field extension $l$ of degree two if $d\equiv 1\mod 4$.
@@
Since $X$ is generic in $\mathcal{L}_{d}$, the elliptic fibration $\pi$ is not isotrivial. Therefore, the base change by a multisection is a non-trivial elliptic fibration with an elliptic base and one can apply Silverman’s Specialization theorem ([Sil83, Theorem C]) to the base change. In particular, the fibers of $\pi$ above the points $t$ in the image of $\pi|_{B}$ verify $r_{t}>r$.
@@
If $d\geq 4$ is even then $X$ admits a realization as a double cover of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ branched over a curve of bidegree $(4,4)$. In this case, the genus 1 fibration is given by the pull-back of the pencil of curves of bidegree $(0,1)$ (resp. $(1,0)$) and there is a pencil of bisections, of which $B$ is a member, given as the pull-back of the curves of bidegree $(1,0)$, (resp. $(0,1)$). Both pencils are clearly defined over the same field $k$. Moreover, the generic fibers of the two genus 1 pencils intersect in two geometric points that are (potentially) conjugate under $\operatorname{Gal}(\bar{k}(t)/k(t))$ and defined over an extension $l(t)/k(t)$ of degree at most 2. In particular, $B$ intersects each fiber $\pi^{-1}(t)$ in two points $\operatorname{Gal}(\bar{k}(t)/k(t))$-conjugate points $p(t),p^{\prime}(t)$. Fix $l=l(t)$, for a $t$ such that $p(t)$ is of infinite order in the Mordell–Weil group of $B$. As the bisections belong to an infinite family, we can conclude that infinitely of them are elliptic curves with a positive rank over $l$, proving the claim. ∎

### v2

@@ 3.2.1. Automorphisms of K3 surfaces in $\mathcal{L}_{d}$
Let $X$ be a general K3 surface in a family $\mathcal{L}_{d}$. Fix the base $\{u_{1},u_{2},b_{3}\}$ of $\Lambda_{d}$ considered in proof of Lemma 3.8. The elliptic involution on each fiber of $\pi$ extends to an automorphism on $X$ that acts on the chosen base of $\Lambda_{d}$ as follows:
@@
If $d\geq 2$, the translation by the section $S_{1}$ is an automorphism of infinite order of the K3 surface, which acts on the base $\{F,\mathcal{O},b_{3}\}$ as follows:
@@
If $d=1$, the previous matrix is still associated with an isometry of the lattice $\Lambda_{d}$, but this isometry cannot be associated with an automorphism since it maps $F-b_{3}$ to $F-b_{3}$. The former is an effective class corresponding to the component of $I_{2}$-fiber meeting the zero section. The latter is not effective since $(-F-b_{3})\mathcal{O}=-1$.
@@
If $X$ is general in its family, then it is $\mathrm{Aut}$-general, i.e., its automorphism group contains only symplectic automorphisms of infinite order (e.g. $T_{1}$) and non-symplectic automorphisms of order 2 (e.g. $\varepsilon$). In particular, this implies that $X$ does not admit isotrivial elliptic fibrations with $j$-invariant of the generic fiber equal to 0 or $1728$. Indeed, otherwise the complex multiplication on each fiber would induce an automorphism on the surface which is non–symplectic and has order 3 or 4, respectively. If an elliptic fibration is isotrivial with a constant $j$-invariant not equal to 0 or 1728, then the reducible fibers are forced to be of type $I_{0}^{*}$, and in particular with non reduced components. By Lemma 2.13, these isotrivial fibrations cannot appear if the Picard number of $X$ is at most 3, i.e. for the surfaces considered in this paper.
@@ 3.3. Smooth rational bisections on elliptic K3 surfaces
Let $X$ be a general K3 surface in $\mathcal{F}_{\rm{rat}}^{x}$ and recall that $x\geq 0$ since it is the intersection number of two irreducible distinct curves. The classes $F$, of the fiber of the fibration, $\mathcal{O}$, of the zero section of the fibration, and $B$, of the rational bisection, are contained in the Picard lattice of $X$. Let us denote $\Phi_{x}:=\left[\begin{array}[]{rrr}0&1&2\\ 1&-2&x\\ 2&x&-2\end{array}\right]$ the lattice $\langle F,\mathcal{O},B\rangle$. By the base change $\{F,\mathcal{O},(-4-x)F-2\mathcal{O}+B\}$, we obtain a lattice isometric to $\Phi_{x}$ which is necessarily primitively embedded $\operatorname{Pic}(X)$. Indeed, we already observed that the sublattice $\langle F,\mathcal{O}\rangle$ has to be primitively embedded in $\operatorname{Pic}(X)$; the class $(-4-x)F-2\mathcal{O}+B$ is not a divisible class in $\operatorname{Pic}(X)$, since the coefficient of $B$ is 1; there are no even overlattices of finite index of $\langle F,\mathcal{O},(-4-x)F-2\mathcal{O}+B\rangle$ in which both $\langle F,\mathcal{O}\rangle$ and $\langle(-4-x)F-2\mathcal{O}+B\rangle$ are primitively embedded. Hence, we proved that the family $\mathcal{F}_{\rm{rat}}^{x}$ is the family of the $\Phi_{x}$-polarized K3 surfaces.
To compare the families $\mathcal{F}_{\rm{rat}}^{x}$ and $\mathcal{L}_{2x+5}$, it suffices to show that the lattices $\Phi_{x}$ and $\Lambda_{2x+5}$ are isometric. We consider again the base $\{F,F+\mathcal{O},(-4-x)F-2\mathcal{O}+B\}$ of $\Phi_{x}$. The intersection form on this base is $U\oplus\langle-2(2x+5)\rangle$. This proves that $\operatorname{Pic}(X)$ is isometric to $\Lambda_{d}$ if $d=2x+5$, with $x\geq 0$.∎
@@
If $\operatorname{NS}(X)\simeq\Lambda_{3}$ (i.e. $d=3$), then the class $B=3F+2\mathcal{O}+b_{3}$ as in (3.2) is still a (2)(-2)-class with intersection 2 with $F$, but it does not correspond to an irreducible curve, since $B\mathcal{O}=-1<0$ and $\mathcal{O}$ is an irreducible curve. Indeed, the class $B$ is not an irreducible bisection, but the sum of two sections, namely $S_{1}=3F+\mathcal{O}+b_{3}$ and $\mathcal{O}$.
@@
The cover involution of $f:X\rightarrow\mathbb{P}^{2}$ is the involution $T_{1}\circ\varepsilon$ and preserves the class of the bisection $B$.
@@ 3.4. Smooth genus 1 bisections on elliptic K3 surfaces
[…]
@@
The base change between $\{F,\mathcal{O},B\}$ and $\{u_{1},u_{2},b_{3}\}$ gives $B=b_{3}+(4+x)F+2\mathcal{O}=(2+x)u_{1}+2u_{2}+b_{3}$, and this allows one to compute the image of $B$ for the automorphisms $T_{1}$, $\varepsilon$ and their composition:
@@
The cover involution of $f:X\rightarrow\mathbb{P}^{1}\times\mathbb{P}^{1}$ is the involution $T_{1}\circ\varepsilon$. It preserves the class of the bisection $B$.
@@ 4. Higher degree multisections
We consider here multisections of degree higher than 2. The main ideas and the proofs are very similar to those in the previous section, where we considered multisections of degree 2. Indeed, as in that case, we discuss the 18-dimensional (resp. 17-dimensional) families of K3 surfaces with a genus 1 fibration (resp. elliptic fibration) admitting a smooth rational or genus 1 multisection of degree $m>2$. Moreover, for low values of $m$, namely $m=3,4,5$, one is also able to describe projective models of the generic K3 surface in the constructed families.
@@
Let $X$ be a K3 surface with a genus 1 fibration $\pi$ and $\rho(X)=2$. Then $\operatorname{Pic}(X)\simeq\Gamma_{m,c}:=\left[\begin{array}[]{cc}0&m\\ m&2c\end{array}\right]$ where $m$ is the minimal degree of a multisection of $\pi$ and $1\leq c\leq m-2$. Moreover, the following hold.
  • i)
    The degree of any multisection of $\pi$ is a multiple of $m$.
  • ii)
    There is a smooth rational multisection of degree $d$ of $\pi$ if and only if $d=m$, where $m$ is the minimal degree of a multisection of $\pi$, and $\operatorname{Pic}(X)\simeq\Gamma_{m,-1}$.
  • iii)
    If $d=gcd(m,c)$ and $m=dm^{\prime}$, then there there is a smooth genus 1 multisection of degree $mm^{\prime}$.
Proof.
The class of the fiber of $\pi$, $F$, is primitive in $\operatorname{Pic}(X)$, and hence there exists a basis of $\operatorname{Pic}(X)$ which is given by $\{F,M\}$ with $FM=m>0$. In particular, chosen $M$ to be the class of an irreducible curve, $M$ is a multisection of degree $m$ of $\pi$. The intersection form computed on these classes is $\Gamma_{m,c}$ and the condition on $c$ up to isometries of the lattices is given in [vG05, Proposition 3.7]. Each class in $\operatorname{Pic}(X)$ can be written as $D=\alpha F+\beta M$, $\alpha,\beta\in\mathbb{Z}$. Therefore, $FD=m\beta$, which yields i). If $D$ is an irreducible curve, and in particular a multisection, then $FD=m\beta>0$ and it is $m$ if and only if $\beta=1$. The self-intersection of $D$ is $2\beta(\alpha m+c\beta)$. If $D$ is the class of a smooth rational multisection, then $\beta(\alpha m+c\beta)=-1$, which implies $\beta=1$ and $c=-1-\alpha m$. In particular, $D=\alpha F+M$, which is not irreducible unless $\alpha=0$. The irreducible rational multisection is then obtained for $\alpha=0$ and so c=1c=-1. So $D$ coincides with $M$ and $D$ has degree $FM=m$. Viceversa, if $\operatorname{Pic}(X)\simeq\Gamma_{m,-1}$, then $M$ (the second generator of the basis of $\operatorname{Pic}(X)$) is a rational multisection. This concludes the proof of ii).
Finally, for iii), if $D$ is the class of a smooth genus 1 multisection, then $\alpha m+c\beta=0$. In particular, if $c=dc^{\prime}$, the class $c^{\prime}F+m^{\prime}M$ has self intersection 0 and is primitive. Hence, it corresponds to the fiber of a genus 1 fibration, whose smooth fibers are infinitely many genus 1 multisections of $\pi$, of degree $(-c^{\prime}F+m^{\prime}M)F=m^{\prime}m$. ∎
We observe that if $m=3$, i.e. $X$ is a K3 surface admitting a genus 1 fibration with a trisection, $X$ can be polarized with three non-isometric lattices that are $\left[\begin{array}[]{cc}0&3\\ 3&2c\end{array}\right]$, with c=1,0,1c=-1,0,1. They correspond to the cases in which the trisection $M$ in the previous proof has genus 0,1 or 2, respectively. When $m$ grows, more possibilities for the polarizing lattices appear, see [vG05, Proposition 3.7].
For low values of $m$, we give nice geometric models of the K3 surfaces and the genus 1 fibration (see [vG05, Paragraph 5.2]): let $M$ be a smooth rational multisection of degree $m$ of the genus 1 fibration whose class of the fiber is $F$, then $H:=F+M$ is an ample divisor, $\varphi_{|H|}(X)\subset\mathbb{P}^{m}$ and there is a pencil of hyperplanes in $\mathbb{P}^{m}$ containing the degree m2m-2 curve $C:=\varphi_{|H|}(M)$.
@@
In all the cases above, the singular fibers of the fibration correspond to hyperplanes which cut on $\varphi_{|H|}(X)$ the curve $C$ and a singular curve. Since $M$ is a multisection of the fibration, we have a map $M\rightarrow\mathbb{P}^{1}$ to the base of the fibration. It is branched over fibers cut out by hyperplanes whose intersection with $\varphi_{|H|}(X)$ consists of $C=\varphi_{|H|}(M)$ and a curve which intersects $C$ in at least one point with multiplicity higher than 1. For a general K3 surface, the hyperplanes corresponding to singular fibers and the ones corresponding to branch fibers for $M\rightarrow\mathbb{P}^{1}$ do not coincide and hence the multisection $M$ is generically a saliently ramified multisection.
@@
Similarly to the proofs of Proposition 3.13 and Lemma 3.19, it suffices to look for an isometry between the lattice $\langle F,\mathcal{O},M\rangle$ and a lattice $\Lambda_{d}$ for a certain integer $d$, where $M$ is the class of an irreducible curve with $FM=m$, $\mathcal{O}M=x\geq 0$ and $M^{2}=-2$, or $M^{2}=0$ accordingly to the genus of the multisection. More explicitly, if $M^{2}=-2$, then $\langle F,\mathcal{O},M\rangle\simeq\langle F,\mathcal{O},(-2m-x)F-m\mathcal{O}+M\rangle\simeq U\oplus\langle-2(m^{2}+mx+1)\simeq\Lambda_{(m^{2}+mx+1)}$, which does not admit even finite index overlattices in which both $\langle F,\mathcal{O}\rangle$ and $\langle(-2m-x)F-m\mathcal{O}+M\rangle$ are primitively embedded. This implies that $NS(X)\simeq\Lambda_{(m^{2}+mx+1)}$, which forces $d\equiv 1\mod m$ and $d\geq m^{2}+1$ since $x\geq 0$. The argument is analogous if $M^{2}=0$, by observing that $\langle F,\mathcal{O},M\rangle\simeq\langle F,\mathcal{O},(-2m-x)F-2\mathcal{O}+M\rangle\simeq\Lambda_{(m^{2}+mx)}$.∎
@@
In Theorem 3.2, we find smooth bisections of low genus (0 or 1) on the K3 surfaces in all the families $\mathcal{L}_{d}$ with $d\geq 4$; the smooth multisections of low genus and degree $m>2$ appear for higher values of $d$. So we do not find smooth saliently ramified multisections of degree $m>2$ on elliptic fibrations that do not admit smooth saliently ramified bisections. We obtain the following result.
@@ 5. Rank jumps for K3 surfaces in $\mathcal{L}_{d}$
Therefore, the only statement which has not yet been proved concerns the field of definition $k$ (resp. $l$), namely, to show that the sets $\mathcal{R}(X,\pi,k)$ (resp. $\mathcal{R}(X,\pi,l)$) are infinite if $d\equiv 3\mod 4$ (resp. $d\equiv 1\mod 4$). In particular will verify that $B$ is defined over $k$ if $d$ is odd, that it has infinitely many $k$-points if $d\equiv 3\mod 4$ and that it has infinitely many $l$-points for a certain field extension $l$ of degree two if $d\equiv 1\mod 4$.
@@
Since $X$ is generic in $\mathcal{L}_{d}$, the elliptic fibration $\pi$ is not isotrivial (see Section 2.3.3). Therefore, the base change by a multisection is a non-trivial elliptic fibration with an elliptic base and one can apply Silverman’s Specialization theorem ([Sil83, Theorem C]) to the base change. In particular, the fibers of $\pi$ above the points $t$ in the image of $\pi|_{B}$ verify $r_{t}>r$.
@@
If $d\geq 4$ is even then $X$ admits a realization as a double cover of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ branched over a curve of bidegree $(4,4)$. In this case, the genus 1 fibration is given by the pull-back of the pencil of curves of bidegree $(0,1)$ (resp. $(1,0)$) and there is a pencil of bisections, of which $B$ is a member, given as the pull-back of the curves of bidegree $(1,0)$, (resp. $(0,1)$). Both pencils are clearly defined over the same field $k$. Moreover, the generic fibers of the two genus 1 pencils intersect in two geometric points that are (potentially) conjugate under $\operatorname{Gal}(\bar{k}(t)/k(t))$ and defined over an extension $l(t)/k(t)$ of degree at most 2. In particular, $B$ intersects each fiber $\pi^{-1}(t)$ in two $\operatorname{Gal}(\bar{k}(t)/k(t))$-conjugate points $p(t),p^{\prime}(t)$. Fix $l=l(t)$, for a $t$ such that $p(t)$ is of infinite order in the Mordell–Weil group of $B$. As the bisections belong to an infinite family, we can conclude that infinitely of them are elliptic curves with a positive rank over $l$, proving the claim. ∎